A new business opportunity has a 85% chance of being worth $750000 next year and a 15% chance of being worth $150000. The appropriate expected rate of return is 11%.     Assume that the new opportunity will be financed with a$540000 bank loan. What rate of return will the levered equity holder earn if the business pays off $750000 in one year?

A. 30.59%
B. 26.14%
C. 336.23%
D. 951.98%
Answer #1

Current or Expected Value of Equity=(0.85*MAX(750000-540000,0)+0.15*MAX(150000-540000,0))/1.11=160810.8108

Return to levered equity=MAX(750000-540000,0)/160810.8108-1=30.59%